Solve
2/z = 3/z - 1/11
Factor Completely
4r^2-81
Factor the trinomial
c^3-2c^2-63c
Use the Foil method to find the product
(7x^4-2)(x^7-4)=
Multiply
(a+d)(a^2-ad+d^2)=

Answer by shree840(260)

There are 5 problems
1.
2/z=3/z-1/11 multiply by z
2=3-z/11
z/11=3-2=1
z=11
2.4r^2-81 this is a^2-b^2 form =(a+b)(a-b)
(2r+9)(2r-9)
3.c^3-2c^2-63c
take c common
c(c^2-2c-63)
c(c^2-9c+7c-63=
c(c-9)(c+7)
4.(7x^4-2)(x^7-4)
foil stands for first outer inner last
7x^4*x^7-7x^4*4-2x^7+(-2)(-4)
=7x^4-28x^4-2x^7+8
=-2x^7-21x^4+8
5.(a+d)(a^2-ad+d^2)
we will still use foil but we will multiply first with a all the numbers in the second bracket and then we multiply all the numbers in the second bracket with d
a^3-a^2d+ad^2+a^2d-ad^2+d^3
=a^3+d^3 {everything else cancels}